OpenStudy (anonymous):

is the vector <1,-3,2> orthogonal (forms a 90 degree angle) with the line given by the following parametric equations: x=1+t, y=-2+2t,z=5

OpenStudy (anonymous):

How can you find if two things are orthogonal? Pro tip: Its a type of product.

OpenStudy (amistre64):

the vector of the line is the "t" coeffs. <1,2,5>; now dot it to the suspected vector and see if we get the 0

OpenStudy (amistre64):

ack ... that 5 got me lol there is no t in z

OpenStudy (amistre64):

<1,2,0>

OpenStudy (amistre64):

i get 1-6 ... so no zero

OpenStudy (anonymous):

Ah ok, thanks mate

OpenStudy (across):

You have a vector \[\vec{a}=\left \langle 1,-3,2 \right \rangle\]And the parametric equations \[x=1+t\]\[y=-2+2t\]\[z=5\]These equations can be de-parametrized as \[\vec{r}(t)=(1,-2,5)+t\left \langle 1,2,0 \right \rangle\]Let's set \[\vec{b}=\left \langle 1,2,0 \right \rangle\]Now, \[\vec{a}\cdot\vec{b}=\left \langle 1,-3,2 \right \rangle \cdot\left \langle 1,2,0 \right \rangle=1-6+0=-5\]Therefore, they're not orthogonal.